OpenStudy (anonymous):

Write an equivalent equation of 12 - 3x = y Solve the s ystems of equations using the elimination method. x + 3 = 7 and 3x +4y = 7

OpenStudy (anonymous):

Please show your process if you can :)

OpenStudy (anonymous):

No i don't

OpenStudy (anonymous):

I know that they equal to eachother and that they are on the same line on a graph.

OpenStudy (skullpatrol):

The first step in learning algebra is to become familiar with its vocabulary. Equivalent equations are equations that have the same solution set over a given domain.

OpenStudy (anonymous):

Solution Set over a given domain?

OpenStudy (skullpatrol):

So to write an equivalent equation of 12 - 3x = y you just rewrite it in an equivalent form.

OpenStudy (skullpatrol):

Use the idea of balance to solve for x.

OpenStudy (anonymous):

Re-writing in equivalent form... Sounds pretty confusing.

OpenStudy (anonymous):

How would i start?

OpenStudy (skullpatrol):

$$\Huge 12 - 3x = y$$ $$\Huge 12 - 3x +3x= y+3x$$ $$\Huge 12 = y+3x$$ $$\Huge 12 -y= y+3x-y$$ $$\Huge 12 -y = 3x$$ $$\Huge \frac{1}{3}(12 -y) =\frac{1}{3} 3x$$ $$\Huge 4 -\frac{1}{3}y = x$$

OpenStudy (skullpatrol):

$$\Huge x = 4 - \frac{1}{3}y$$$$\Huge\text{and}$$ $$\Huge y=12 - 3x $$ $$\Huge\text{are equivalent equations.}$$

OpenStudy (anonymous):

How did you get from 1/3(12-y) = 1/3 3x to 4-1/3y = x?

OpenStudy (skullpatrol):

the distributive property a(b+c)=ab+ac

OpenStudy (anonymous):

How did the 12 become 4 ?

OpenStudy (skullpatrol):

(1/3)(12) = 4

OpenStudy (anonymous):

Oh i see.

OpenStudy (anonymous):

Do you always have to flip the variables to the opposite sides when rewriting into an equivalent equation?

OpenStudy (skullpatrol):

No, I did it for appearance.

OpenStudy (anonymous):

So when writing Equivalent Equations.. lets see if i made one up 4 -4x = y Would it be like +4x 4 = y + 4x -y 4 -y = 4x (1/4)(4/1) 4 -y = x?

OpenStudy (anonymous):

And what about my last question about the elimination method.

OpenStudy (skullpatrol):

Oops, your last step should be multiply both sides by 1/4

OpenStudy (anonymous):

MMM isn't that what i did?

OpenStudy (anonymous):

1/4)(4/1) 1 x 4 = 4 1 x 4 = 4

OpenStudy (anonymous):

oh it should have been 1?

OpenStudy (anonymous):

1 -y = x?

OpenStudy (skullpatrol):

4 -y = 4x (1/4) (1/4)

OpenStudy (skullpatrol):

4 -y = 4x (1/4)(4-y) = (1/4)4x

OpenStudy (anonymous):

Oh we have to do it to the entire side.

OpenStudy (anonymous):

1 -1/4y = x?

OpenStudy (skullpatrol):

Yes :-)

OpenStudy (anonymous):

Oh i see, thank you very much. :) And how would i solve the system of equations using the elimination method if you don't mind me asking.

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

MMmi know when you use the elimination method you're supposed to get rid of some variables maybe by using their opposite. But i wouldn't know where to start.

OpenStudy (skullpatrol):

Let's look at x + 3 = 7 and 3x +4y = 7 and decide what is a convenient variable to solve for? We have two choices, right?

OpenStudy (anonymous):

Yeah x and y

OpenStudy (anonymous):

3 didn't have a variable which confused me.

OpenStudy (anonymous):

We would probably solve for x?

OpenStudy (skullpatrol):

yes

OpenStudy (skullpatrol):

which equation do you want to solve for x

OpenStudy (anonymous):

x +3 = 7

OpenStudy (skullpatrol):

ok, do it please

OpenStudy (anonymous):